OpenStudy (anonymous):

Find a third-degree polynomial equation with rational coefficients that has the roots 2 and -3+i

OpenStudy (athe):

\[(x-2)(x^2+9x+10)=0 \rightarrow x^3+7x^2-8x-20=0\] recomendation \[(x-2)(x-(-3-i))(x-(-3+i))=0\]